The product of two consecutive odd integers is 143. Find the integers
Answer by amarjeeth123(569)

Let the consecutive odd integers be x and (x+2)
x(x+2)=143
x^2+2x-143=0
x^2+13x-11x-143=0
x(x+13)-11(x+13)=0
(x-11)(x+13)=0
x=11 is the only solution.
The two consecutive odd integers are 11 and 13.